OpenStudy (anonymous):

the time it takes to set up tables for a banquet dinner varies inversley as the # of people setting tables. it takes 2 hours for 4 people to set up tables . how long will it take six people to do the job

OpenStudy (anonymous):

t= time to set up n= number of people inverse proportional follows this equation: k=n*t where k is a constant first solve for k 4*2=k k=8 so 8=n*t 8=6*t t=8/6

OpenStudy (anonymous):

thanxx